OpenStudy (anonymous):

The product of 8/15 6/5 1/3 is A)48/15 B)16/75 C)48/30 D)16/15

OpenStudy (anonymous):

B. 6/5 x 1/3 = 2/5 2/5 x 8/15 = 16/75